OpenStudy (egbeach):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) =

geerky42 (geerky42):

4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) \(\textbf=\) ??? what?

OpenStudy (egbeach):

it doesnt say... thats how it was written

OpenStudy (zzr0ck3r):

ouch

OpenStudy (zzr0ck3r):

it is not a statement...

geerky42 (geerky42):

we cannot prove expression "4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2)"

geerky42 (geerky42):

I mean it doesn't make sense.

OpenStudy (egbeach):

thats what i thought

OpenStudy (zzr0ck3r):

I cant find the formula elsewhere.

geerky42 (geerky42):

What I realize though is that pattern is supposed to be \(n(n+2)\) starting at \(n=4\), not \(4n(4n+2)\) Because otherwise pattern is supposed to be \(4\cdot6+8\cdot10+12\cdot14+\cdots\)

OpenStudy (egbeach):

okay i found it. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = (4(4n+1)(8n+7))/6

OpenStudy (zzr0ck3r):

right @geerky42

OpenStudy (zzr0ck3r):

Do you want to see if it matches the formula she found?

geerky42 (geerky42):

Maybe I misinterpreted the pattern, it's just \(\displaystyle \sum_{i=4}^{4n}i(i+2)\) ?

OpenStudy (egbeach):

wait i found the answer since the statement is false. thank you though